OpenStudy (anonymous):

what is the answer of;4×4+4×4+4-4×4=?

OpenStudy (rane):

thats basics, something that u have to and must know

OpenStudy (kc_kennylau):

http://bit.ly/1cTi0Q0

OpenStudy (rina.r):

cool their u go @kc_kennylau answered ur question..

OpenStudy (kc_kennylau):

xD

OpenStudy (skullpatrol):

google is not a way to learn math

OpenStudy (kc_kennylau):

true there

OpenStudy (anonymous):

4*4+4*4+4-4*4=16+16+0*4=16+16+0=32

OpenStudy (kc_kennylau):

It's 4-(4*4) not (4-4)*4 :)

OpenStudy (anonymous):

then the answer is 20

OpenStudy (anonymous):

its 20

OpenStudy (anonymous):

16+16+4-16=20

OpenStudy (kc_kennylau):

PEMDAS: P stands for Parentheses, always do the things inside parentheses first :) E stands for Exponents, you'll learn it later :) MD stands for Multiplication and Division, do whichever comes first :) AS stands for Addition and Subtraction, do whichever comes first :) \[\large\begin{array}{llrl}&4\color{green}\times4\color{blue}+4\color{green}\times4\color{blue}+4\color{red}-4\color{green}\times4\\=&4\color{green}\times4\color{blue}+4\color{green}\times4\color{blue}+4\color{red}-4\color{green}\times4&\mbox{step one: }&\mbox{parentheses}\\=&4\color{green}\times4\color{blue}+4\color{green}\times4\color{blue}+4\color{red}-4\color{green}\times4&\mbox{step two: }&\mbox{exponents}\\=&16\color{blue}+16\color{blue}+4\color{red}-16&\mbox{step three: }&\mbox{multiplication and division}\\=&20&\mbox{step four: }&\mbox{addition and subtraction}\end{array}\]

OpenStudy (anonymous):

by BODMAS rule, (4*4)+(4*4)+4-(4*4)=16+16+4-16=36-16=20